The optical table with the filter cavity remains very clean: 0/0/0 over a 60 second measure.
[Daniel, Torrey]
We have a rough round trip of the cavity aligned. Added a halfwave plate, PBS, and PD as an input power reference. We made a mode matching solution but found that the few inches before the cavity are not allowable to use as a shifting range for the lens as the beam from the round trip of the cavity will not clear the lens. Quick trig says we can't put a lens within 7 inches of the input of the cavity. Will redo this after lunch.
With permission, we borrowed a beam profiler from the QIL Lab. Attached is an image of the two profilers in the box. This loan is also noted in this logpost on the QIL Elog
I 3D printed an aligment tool that goes inside the bowtie cavities to aid alignment. The beam is suppose to go through the two holes. The hole diameter is 0.25 in so the beam shouldn't clip, although we should probably remove this during actual operation since there are other sites in which the beam almost clips.
Attached is the 3D printed part as a SolidWorks and STL file.
I 3D printed a holder for the Beckman Coulter HHPC 3+ particle counter so that it can charge and be run simultaneously. This holder screws into the particle counter and into the optics table so that it doesn't tip over. There is also a taller prototype version which doesn't screw into the table in case wires that plug into the particle counter are too tall and stiff.
Attached is the good version of this 3D print as a SolidWorks and STL file and the previous version only as an STL file. I did not recreate the particle counter in SolidWorks. If I were to 3D print this again, I would increase the internal radius on the design to create a more snug fit with the particle counter. This is because the particle counter only has 1 available hole to use. Note: supports were used in the 3D print.
I tested the LBP2-HR-VIS2, which I borrowed from Nick Hutzler's lab, as a beam propfiler. It is speced to 190 nm - 1100 nm, but I wanted to see if it worked for 1550 nm light. As an initial test, I put it in front of ~2 mW of 775 nm light and it profiled it well. Unfortunately, it couldn't see any of the ~10 mW 1550 nm light even with all of the ND filters off. A footnote says, "Although our silicon cameras have shown responses out to 1320 nm, it can cause significant blooming which could lead to significant errors of beam width measurement." Therefore, I don't think this product will work as a beam profiler replacement.
I added beam dumps to the SHG farm. The outputs were previously unblocked and going off the table. I also added a beam block by the filter cavity so that it is not nudged.
[Daniel,Torrey]
Construction of the first filter cavity has begun. We mounted all the required optics leading up to the cavity this morning. Now we have light from the power distribution center in a 5m fiber piped over to the other end of the table, going through both EOMs, and into the filter cavity. We are ready for the good mirrors now.
[Sander,Daniel,Ian,Torrey]
We moved the sled with the filter cavity to the cleaner table (table directly to the left when entering the room) and installed 3 newport 1550 laser line mirrors at each point on the filter cavity except for the input. We used the whole punched IR card to allow the input beam to enter and aligned the reflection of the 3 mirrors, i.e. one round trip in the cavity, to match the input beam. The beam incident on the cavity is now from the path of the power distrubution center previously used for Daniel's crusher tests.
[Torrey, Daniel]
Within the Michelson Interferometer to test the piezo, we disassembled the Piezo assembly. On the back of mirror, there appears to be an etching (see attached photo).
We then added a 1/16 in thick steel spacer with a 1 in OD and 0.25 in ID between the mirror and piezo (see attached photo).
We slightly tweeked the alignment to get good contrast and took open and closed loop transfer functions (see attached photo and data).
We expect the fundamental resonance to be changed by the following: \[f_2 = f_1 \sqrt{\frac{m_1}{m_2}} \] , where f_1 = 8.5 kHz , m_1 = 6.7 g (the mass of the mirror), and m_2 = 11.6 g (the mass of the mirror plus the spacer).
This predicts f_2 = 6.6 kHz , but we measure f_2 =6 kHz .
Even though there is this 10% discrepancy, this gives us some confidence we are looking at the first longitudinal eigenmode at 8.5 kHz without the spacer.
[Torrey]
Broke out the particle counter. Few readings around the room. All readings are for particle sizes .3 um/.5 um / 1 um.
1) Inside clean area near IFO test set up:111/27/11
2) Outside clean area near NUK1: 7303/1459/472
3) Inside clean area near SHG: 543/105/32
4) Inside clean area, middle of table nearest the door: 3/0/0
I think only one fan is turned on in the clean area with the SHG/power distribution/etc on it. Also kind of silly, you can't stand the counter up while its charging... Have it charging on the racks above the second table in the clean area at the moment.
[Alex, Boris]
Boris and I ran the first GQuEST SNSPD tests this past friday!!
We ran IV curve traces to determine the switching current of the SNSPD to be about 21 uA (image 1).
We have a nice image of the pulse seen on the oscillascope (image 2).o
Next we ran a count rate test and a dark count rate test. The two tests together gave us our output efficiency which was in the range of 85-90% currently. (image 3&4) where the input number of photons was ~29,000.
We then ran an hour long test for the dark count with the fiber input connector terminated which I will post a more in depth graph for tomorrow.
All seems well for the GQuEST SNSPD at this time, and we will next be opening the fridge this coming wednesday and replacing the SNSPD case for the completely dark case and rerunning our DCR tests.
All data is in the next cloud: Nextcloud\GQuEST\SNSPD\Efficiency_testing\DATA\11102023
[Ian, Torrey, Sander]
We cleaned the lab of all of the trash we could find. We also unpacked a number of things including the LIOP-TEC Mirror mounts. They are now all put away and we have places for all of them. We organized the large cabinet with the drawers unpackaging everything we could and making sure that is was ready for dividers. We should do this as things arrive so that we don't need to unpackage mounts and such as we need them. We will organize everything better when the dividers for the drawers arrive hopefully by the 19th.
We were able to remove more than 5 trashbags worth of trash from the lab and we aren't done. We should make sure that it does not get this bad again. We should also buy more trash cans so that they are better placed when we have trash. I would suggest one by the cabinet with the drawers.
[Alex, Ian, Torrey]
B150 has a bunch of old AOMs of various shapes and sizes. We set up a circuit to test them. We used the signal analyser to output a signal into a mixer, have that signal go to the AOM and reflect back to the input of the analyser. The idea being we should see dips in this signal near the AOMs resonance. We did this for ~8 AOMs that we had, and have identified ones that are clearly not in working condition. The remaining ones in the ULine cabinet should all be working, with resonances observed between 60-150MHz. We will need a beam profiler to continue with AOM testing from here (maybe one of the NIR cameras that are suppose to arrive this week).
While testing the crusher, the beam profiler (NS2s-GE/5/5-STP Model PH00460) fell off the table as I was adjusting it. It turns on, but it does not detect a beam when it is in front of it. When it is on, the scanning slit makes a rubbing sound. I reached out to Ophir to fix it.
[Daniel, Torrey]
The profiler was fixed by Ophir and shipped back. It was mounted using the wider BA2 and used without issue.
[Daniel, Sander]
Using the Moku:Pro, we measured the (open-loop) transfer function of the mounted ring piezo mirror actuator in a Michelson interferometer. The piezo-mirror assembly seems to have a fundamental resonance at around 8.7 kHz.
We sent a swept sine signal from the Moku to the piezo-mirror assembly (through the Thorlabs piezo controller). The actuated mirror is an end mirror of a Michelson interferometer, and we read out the response from the signal of a photodiode at the Michelson output. We used the Moku:Pro in multi-instrument mode, with the following instruments (see screenshot):
Slot 2: Laser Lockbox to perform a fringe lock.
Slot 3: Frequency Response Analyser to send a swept sine, record the response, and compute the transfer function.
Slot 4: PID controller to add the control signal and swept sine signal into one output signal.
We performed the transfer function measurement in two configurations:
I) Without feedback control of the Michelson fringe operating point (loop OPEN)
For this we first put the Michelson at roughly mid-fringe, then disabled the laser lockbox digital output and opened the loop. The transfer function is computed for the ratio of the Moku output signal (containing only the swept sine) and Moku input signal (containing the PD signal). See attached screenshot.
II) With feedback control of the Michelson operating point (loop CLOSED)
For this the control loop is closed and the Michelson is maintained at the mid-fringe. The transfer function is computed for the ratio of the Moku output signal (containing the swept sine and the control signal) and Moku input signal (containing the PD signal). See attached screenshot.
Both measured transfer functions appear very similar, as they should in theory. It appears the piezo-mirror assembly goes through resonance at roughly 8.7 kHz, although the fundamental peak is low-Q and hardly visible.
[Ian, Torrey]
Ian and I wanted to confirm these measurements. Here is our data and a screenshot of the graphs. We made minor tweaks to the amplitude of the signal going to the piezo and the loops in the controller. This yielded essentially the same results as above.
This is a data dump for the random assortment of AOM's and RF drivers we have in from the previous lab.
Power sources:
There is one of these. Should be sufficient for intial AOM tests.
AOMS:
1) Brimrose Model TEF-200-50 (this is my best guess, not 100% sure)
(https://www.brimrose.com/free-space-ao/acousto-optic-frequency-shifters).
Model # | Spectral Range (nm) |
Diffraction Efficiency Up To (%) |
Frequency Shift (MHz) |
Active Aperature (mm) |
Material |
---|
TEF-200-50 | 400-1600 | 60 | 200 | 0.75 | TEO2 |
2) NEC A-O Modulator Mod. C8217-A (Discontinued model. No info on these. There are at least 2 of these)
57-78 MHz drive frequency.
3) Isomet AOM Model 1201E
https://isomet.com/PDF%20acousto-optics_modulators/data%20sheets-modnir/M1201-SF40-1p7.pdf
https://neurophysics.ucsd.edu/Manuals/Isomet/AO%20Modulator-Driver%20Series%20230.pdf
FREQUENCY BW POWER
40MHz 20MHz 1.6W
4) There are an additional 3 viable AOMs with not enough distinguishable markings on them to identify what kind they are.
Plan to use the TEF-200-50 to start.
All these AOMS are currently stored in FG6 in the ULine cabinet.
[Ian, Torrey]
We set up one flipper mirror mount as a remote beam block for the power distribution. They can be controlled 3 ways:
1) Manually push the button on the flipper mount (whole point of these is to not have to do this).
2) There is a physical extension of the button (SMA connection) you can push, see attached.
3) There is a USB connection and a little GUI provided by thorlabs. The software is called Kinesis. Attached is what it looks like.
We need a USB hub to do a flipper mirror for all four paths and the SHG temperature controller at the same time.
Retook beam profile coming out of SHG with everything in place. SHG Profile data can be found at Nextcloud/GQuEST/B102/SHG/SHG\ Beam\ 20Profile/SHG_Profile.ipynb. Will use this to match to the fiber inputs at the end of the SHG farm. All 4 paths on the SHG farm should be within an inch of each other so 1 MM solution before the first 50:50 should be sufficient. Will also post a solution for the two lens on either side of the AOM.
The input couplers for fibers we have want ~500 um size beams at the input. The attached photos show an approximate solution in JAMMT, and its experimental implementation. Discrepency arises from the uncertainty of the length in the path from the PBS to the input couplers.
Then I came up with an approximate solution for the lenses before and after the AOM. The full_jammt photo shows this to scale with the space we have on the sled. I think this will work. Ready to fire up an AOM now.
I also noticed a mistake on the SHG sled that had the path lengths differing by a couple inches that has been corrected.
I created a rough MM solution and installed it that puts a waist of ~700 um at M3. Took off M3 and used the profiler to check beam size at this location. My code in finesse says ~605 um is the actual size required, but this will do for now. I also hooked up the EOM and ensured it was working. Installed the curved optic at the M1 location and am ready to try for cavity alignment.